On August 19 you bought the OCT, K=25, call and at the same time you bought the OCT, K=25 put. You hold both options to their expiration. At the options expiration which one will you exercise and what will be your profit/share or loss/share if MMM’s price at expiration were:

5.1 S = 35; 5.2    S = 20;     5.3     S = 25.

NOTE: Your profit is defined as:

The per share cash flow at expiration PLUS the initial CASH FLOW per share

MMM; TUE August 19 2014. St 27.50 CALLS LAST PUTS LAST Sep14 Oct14Jan15Apr15Sep14Oct14Jan15 Apr15 20 25 27.5 30 32.5 35 37.5 8.50 .35 3.50 3.80 .15 .50 1.35 .55 4.20 1.70 .75 35 .24 2.40 2.75 .45 1.30 5.85 8.10 11.0011.74 8.75 .05 1.00 8.85 9.50 .50 .94 12.50
Answer #1

Initial cash outflow per share = Oct K = 25 call premium + Oct K = 25 Put premium = 3.8+0.5 = 4.3

If S = 35

You will exercise the call option since the put option premium would be 0 in this case.

Oct K = 25 call option inflow at expiration = 35-25 = 10

Therefore profit per share = 10 - 4.3 = 5.7

If S = 20

You will exercise the put option since the call option premium would be 0 in this case.

Oct K = 25 put option inflow at expiration = 25-20 = 5

Therefore profit per share = 5 - 4.3 = 0.7

If S = 25

You will not exercise any option because the strike price and price of underlying asset is same, therefore the premium of both the option at expiration would be zero

Therefore Loss per share = 0 - 4.3 = -4.3